### v1

@@ 2.5 Comparative Study
We performed a comparative study with two popular approaches for training the multi-modal LLMs.
1. Simple Adapter with LoRA finetuning [41]; in this approach, one finetunes the language model using parameter-efficient adapters.
2. Bootstrapping Approach [28]; in this approach, only the bridge model is trained while keeping the graph encoder and LLM frozen, which is the method adopted in our work.
Both approaches were evaluated on the same dataset after being trained on identical training sets for the same duration. The results are presented in Table 1. While the Simple Adapter with LoRA finetuning showed improvements over the LLM-only baseline, its performance still lags behind our bootstrapping approach. This suggests that the bootstrapping method is more efficient for this specific problem setup, where the goal is to align material structural data with language representations without extensive finetuning of the underlying models.
| Task | Simple Adapter w/ LoRA Finetune LLM | MatterChat |
| Metallic(Accuracy) | 0.58 | 0.867 |
| Direct Bandgap(Accuracy) | 0.84 | 0.8702 |
| Stability(Accuracy) | 0.741 | 0.8369 |
| Exp Ob(Accuracy) | 0.715 | 0.8119 |
| Is Magnetic(Accuracy) | 0.835 | 0.9343 |
| Magnetic Order(Accuracy) | 0.777 | 0.8477 |
| Formation Energy(RMSE) | 0.212 | 0.1526 |
| Energy Above Hull(RMSE) | 0.1786 | 0.1095 |
| Bandgap(RMSE) | 1.52 | 0.5428 |
@@
| Task | MatterChat | MatterChat w RAG |
| Metallic(Accuracy) | 0.867 | 0.878 |
| Direct Bandgap(Accuracy) | 0.8702 | 0.872 |
| Stability(Accuracy) | 0.8369 | 0.855 |
| Exp Ob(Accuracy) | 0.8119 | 0.857 |
| Is Magnetic(Accuracy) | 0.9343 | 0.9322 |
| Magnetic Order(Accuracy) | 0.8477 | 0.8507 |
| Formation Energy(RMSE) | 0.1526 | 0.1225 |
| Energy Above Hull(RMSE) | 0.1095 | 0.0949 |
| Bandgap(RMSE) | 0.5428 | 0.5070 |
[…]
@@ 3 Discussion
In this study, we present MatterChat, a multi-modal framework that demonstrates superior performance in material property prediction and scientific reasoning tasks compared to traditional models. This performance can be attributed to its ability to integrate structural material data with language-based information, creating a more comprehensive representation of materials. Traditional approaches often rely on text-based representations, such as Crystallographic Information Files (.cif), to encode atomic structures. While .cif files effectively capture detailed structural information, their static format may limit the model’s capacity to learn complex structure-property relationships when used directly as input. In contrast, MatterChat’s multi-modal approach processes structural embeddings derived from atomic graphs alongside textual data, enabling dynamic interactions between atomic configurations and contextual scientific knowledge. This integration enhances the model’s generalization ability across diverse material tasks, as evidenced by its improved accuracy in classification and numerical property prediction tasks.
While MatterChat demonstrates advancements in material property prediction and analysis, several limitations provide avenues for future improvement and exploration.
1. Alignment Limitation:
One key limitation lies in the current level of alignment between the graph-based material(atom) embeddings and the language representations. Although the framework achieves superior performance on trained material property tasks, this behavior-based alignment does not necessarily imply full representation alignment between the two modalities. Representation level alignment remains an area requiring further investigation [42].
2. Data Limitations:
The curated dataset used in this study, while effective for benchmarking, has certain constraints that may impact model generalization ablility. Specifically, the text data for material property tasks consists of a fixed number of paraphrased queries, which could limit the diversity of the language representations. Expanding the dataset to include more comprehensive and diverse textual description scheme, as well as incorporating a broader range of material properties, could address this issue.
3. Frozen LLM:
Using frozen LLMs may limit the framework’s effectiveness in material-related tasks, as they are primarily trained on broad, general-purpose text rather than domain-specific material science literature. Incorporating specialized LLMs trained on material science datasets could potentially enhance performance by improving domain-specific contextual understanding and reasoning.
To address these limitations, future work could focus on:
1. Improving Alignment:
Leveraging advanced techniques to improve representation-level alignment between modalities during end-to-end instructive fine-tuning, such as incorporating contrastive loss at the fine-tuning stage in addition to pretraining to further reinforce structural-textual consistency.
2. Dataset Expansion:
Building a more comprehensive dataset with diverse textual queries and a broader range of material properties to ensure the framework’s applicability to real-world scenarios.
3. Material Generation Capabilities:
Developing a graph generative module to enable the framework to tackle new material design and discovery tasks.
4. Optimizing LLM Integration:
Investigating finetuning strategies for LLMs, such as whether multimodal LoRA finetuning offers advantages over independent LLM finetuning before integration, to improve downstream task performance.
@@ Acknowledgements
Y.T., Z.Y., and A.N. were supported by Laboratory Directed Research and Development (LDRD) funding from Berkeley Lab, provided by the Director, Office of Science, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 and under NERSC GenAI award under No. DDR-ERCAP0030541.

### v2

@@ 2.5 Comparative Study
We conducted a comprehensive comparative study to evaluate the effectiveness of our multi-modal LLM approach against several alternative methods. The summary of inference results on the test set across all material property prediction tasks is presented in Table 1.
The first column shows results from training the multi-modal LLM using a Simple Adapter with Low-Rank Adaptation (LoRA) finetuning[41], where lightweight adapter layers and the LLM (Mistral-7B) are updated during training, following the method in [42].
The second column presents a pure LLM baseline, in which the LLM is finetuned with LoRA on serialized CIF content and user queries treated as tokenized text prompts—without any structural encoding or auxiliary modality.
The third column corresponds to our proposed method based on the Bootstrapping Approach [28], where only the bridge module is trained, while both the graph encoder and the LLM remain frozen. This strategy focuses on aligning material structures with language representations efficiently, avoiding the need for extensive finetuning of the large-scale language model.
All methods were trained on the same dataset under identical conditions and evaluated on the same test splits. The results demonstrate that our multi-modal bootstrapping approach achieves superior performance compared to both baselines. Notably, it does so without modifying the pretrained components, highlighting its efficiency and practicality for structure-aware property prediction tasks using LLM.
We also include results from our approach augmented with multi-modal retrieval-augmented generation (RAG) during inference. This variant further slightly improves inference performance by having taking more similar material samples into consideration during the inference.
| Task | Simple Adapter w LoRA | LoRA LLM only | MatterChat | MatterChat w/ RAG |
| Metallic (Accuracy) | 0.6373 | 0.6864 | 0.8683 | 0.8873 |
| Direct Bandgap (Accuracy) | 0.8629 | 0.7839 | 0.8753 | 0.8797 |
| Stability (Accuracy) | 0.7418 | 0.7944 | 0.8515 | 0.8573 |
| Exp Ob (Accuracy) | 0.7171 | 0.6549 | 0.8504 | 0.8570 |
| Is Magnetic (Accuracy) | 0.8339 | 0.6833 | 0.9368 | 0.9333 |
| Magnetic Order (Accuracy) | 0.7759 | 0.4238 | 0.8570 | 0.8535 |
| Formation Energy (RMSE) | 0.4105 | 1.8059 | 0.1500 | 0.1212 |
| Energy Above Hull (RMSE) | 0.4415 | 0.4051 | 0.1053 | 0.0964 |
| Bandgap (RMSE) | 1.2516 | 1.4725 | 0.5590 | 0.5058 |
@@
Note: RMSE units — Formation Energy and Energy Above Hull in eV/Atom, Bandgap in eV.
@@ 3 Discussion
In this study, we present MatterChat, a multi-modal framework that achieves superior performance in material properties prediction and scientific reasoning tasks by leveraging a more effective representation of materials. A key innovation of MatterChat is its ability to leverage existing advancements in both material science and language modeling by integrating a pretrained material foundation encoder with a pretrained large language model (LLM). Rather than training an entire model from scratch, MatterChat achieves strong performance by training only a lightweight bridge model, efficiently aligning material structure representations with textual understanding while maintaining high accuracy across diverse material science tasks. Moreover, MatterChat is designed for multitask learning, enabling it to handle both classification and numerical property prediction. This capability allows the framework to tackle a diverse range of material science tasks within a unified model. Another advantage of our approach is the use of graph-based structural embeddings instead of relying solely on Crystallographic Information File (.cif) text input. While CIF files encode atomic structures, their text-based format relies entirely on attention mechanisms, which can struggle to explicitly capture geometric symmetries and increase computational overhead due to lengthy tokenization. By processing atomic graphs directly, MatterChat effectively preserves material symmetry and spatial relationships, leading to more accurate structure-property learning while maintaining computational efficiency.
While MatterChat demonstrates strong performance in material property prediction and analysis, several limitations present opportunities for improvement. The alignment between graph-based material embeddings and language representations remains behavior-driven rather than fully representation-level aligned, requiring further investigation. Additionally, the dataset’s fixed paraphrased queries may limit linguistic diversity, suggesting the need for more comprehensive textual descriptions and a broader range of material properties. Using a frozen LLM, primarily trained on general-purpose text, may restrict domain-specific reasoning and can potentially benefit from finetuning on material science datasets and literature. Future work could focus on three key directions to enhance MatterChat. First, improving representation-level alignment between modalities by incorporating contrastive loss during end-to-end instructive fine-tuning, in addition to pretraining, could further reinforce structural-textual consistency. Second, expanding the dataset to improve generalization by increasing linguistic diversity, as the current fixed set of paraphrased queries may limit variability. Incorporating more comprehensive descriptions and a broader range of material properties would enable the model to handle diverse material science tasks more effectively. Finally, to further enhance MatterChat’s generative capabilities, integrating a graph generative module on top of the existing model could enable novel material structure discovery, extending its utility from property prediction to full-scale material design.
@@ 4.3 Embedding Visualization
Data availability
Upon publication, all data that support the plots within this paper and other findings of this study will be available on a public GitHub repository.
Code availability
Upon publication, all codes that support the plots within this paper and other findings of this study will be available on a public GitHub repository.
@@ Acknowledgements
Y.T., Z.Y., and A.N. were supported by Laboratory Directed Research and Development (LDRD) funding from Berkeley Lab, provided by the Director, Office of Science, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. This research used resources of the National Energy Research Scientific Computing Center (NERSC), a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 and under NERSC GenAI award under No. DDR-ERCAP0030541. W.G. acknowledges support from the National Science Foundation through Grant No. 2235276.
Author Contributions Statement
Y.T. and W.G. conceived the idea, with W.X. contributing enhancements and Y.Z. providing additional support. Y.Z. supervised the project. Y.T. constructed overall ML framework and performed ML training/inference. W.X performed the physical model training and visualization. Y.T. and W.X. wrote the manuscript with the help of Y.Z., J.C., J.M., A.N, S.F., B.E. and M.M.